# Prediction of Energy Resolution in the JUNO Experiment

### v1

@@ Abstract
This paper presents the energy resolution study in the JUNO experiment, incorporating the latest knowledge acquired during the detector construction phase. The determination of neutrino mass ordering in JUNO requires an exceptional energy resolution better than 3% at 1 MeV. To achieve this ambitious goal, significant efforts have been undertaken in the design and production of the key components of the JUNO detector. Various factors affecting the detection of inverse beta decay signals have an impact on the energy resolution, extending beyond the statistical fluctuations of the detected number of photons, such as the properties of liquid scintillator, performance of photomultiplier tubes, and the energy reconstruction algorithm. To account for these effects, a full JUNO simulation and reconstruction approach is employed. This enables the modeling of all relevant effects and the evaluation of associated inputs to accurately estimate the energy resolution. The study reveals an energy resolution of 2.95% at 1 MeV. Furthermore, the study assesses the contribution of major effects to the overall energy resolution budget. This analysis serves as a reference for interpreting future measurements of energy resolution during JUNO data taking. Moreover, it provides a guideline in comprehending the energy resolution characteristics of liquid scintillator-based detectors.
@@ 1 Introduction
The discovery of neutrino oscillations has established that at least two neutrino mass eigenstates are massive. This unambiguous conclusion has provided the first evidence of new physics beyond the Standard Model which assumes neutrinos are massless. In the framework of three neutrino oscillations, six fundamental parameters are involved to describe the oscillation phenomena. The solar and reactor neutrino experiments have determined the mass-squared difference $\Delta m^{2}_{21}$ and the mixing angle $\sin^{2}\theta_{12}$. Atmospheric and accelerator experiments have measured $|\Delta m^{2}_{32}|$ (or $|\Delta m^{2}_{31}|$) and $\sin^{2}\theta_{23}$. Reactor antineutrino experiments have provided the most precise determination of $\sin^{2}\theta_{13}$, and have also achieved a comparable precision of $|\Delta m^{2}_{32}|$ with the measurements from the atmospheric and accelerator experiments. The two independent mass-squared differences and the three mixing angles have been measured with a precision of a few percent [1]. However, the sign of $\Delta m^{2}_{32}$ (referred to as neutrino mass ordering, NMO) and the value of the CP violating phase $\delta_{\textrm{CP}}$ remain unknown, thus, several experiments [2, 3, 4] have been proposed to determine these two crucial parameters.
@@
Over the past few years several factors related to the NMO sensitivity in JUNO have undergone changes. As a result, the JUNO collaboration recently updated the NMO sensitivity using more realistic inputs [14]. These updates include the incorporation of updated reactor thermal powers and new constraints on the reactor neutrino spectrum we expect to obtain from the TAO [15] satellite detector. Additionally, the accidental backgrounds have been revised based on radioassay results of the detector’s raw materials and the estimation of radioactive impurities during the detector installation [16]. The cosmogenic backgrounds have been re-evaluated due to changes in the overburden, and new backgrounds from global reactors and atmospheric neutrinos have been included. Furthermore, the detector energy resolution has been updated through the full detector simulation and event reconstruction, which take into account a better understanding of the detector structure, as well as the performance of the liquid scintillator (LS) and photomultiplier tubes (PMT) [17, 18]. This paper serves as a complementary report to the updated NMO sensitivity paper of JUNO and provides a detailed description of the latest understanding made to the energy resolution prediction in the JUNO experiment, which is essential in the NMO analysis.
@@
In previous LS-based neutrino detectors as the Borexino [19] and KamLAND [20] experiments achieved the energy resolution levels of 5% and 6.5% at 1 MeV, respectively. In these experiments, statistical fluctuations dominate the energy resolution, with the reported PE yields of 511 PE/MeV and 250 PE/MeV, respectively. The JUNO detector was designed with a focus on achieving high LS light yield, high transparency, and high photon collection efficiency to minimize the statistical fluctuation of the detected PE. The LS recipe has been optimized to maximize the light yield using one of the antineutrino detectors in the Daya Bay experiment [21]. Additionally, 17,612 high detection efficiency 20-inch PMTs (LPMTs) and 25,600 3-inch PMTs (SPMTs) have been deployed, resulting in a total photocathode coverage of 78%. However, due to the unprecedented energy resolution requirement of JUNO, several other factors, which are usually negligible in previous experiments, significantly contribute to the energy resolution and must be carefully evaluated. These factors include the quenching effect in LS, Cherenkov radiation, electronics and PMT responses, among others. All of these effects have been modeled in the full detector simulation and will be thoroughly discussed in this paper, along with the impacts from the vertex and energy reconstruction algorithms.
@@
The remaining parts of this article are organized as follows. We first provide an introduction to the JUNO detector in Sec. 2, followed by a summary of the main factors that impact the energy resolution throughout the processes of light production, propagation, and detection in Sec. 3. The full detector simulation is then presented in Sec. 4, focusing on the optical models of the LS and PMTs. Updates and improvements made to these models based on the latest knowledge and measurements are highlighted. The approaches used to determine key parameters in these models are also discussed. Additionally, the electronics simulation is introduced in Sec. 5, which provides a comprehensive model of the detector response. Section 6 describes the calibration and event reconstruction procedures and reports the newly predicted energy resolution in JUNO, providing in Sec. 7 a decomposition of the energy resolution to identify the major contributing factors. Finally, a summary is given in Sec. 8.
@@ 2 The JUNO detector
A schematic drawing of the JUNO detector is shown in Fig. 1. The detector consists of three main components: the central detector (CD), the water pool (WP), and the top tracker (TT).
@@
Scintillation and Cherenkov light produced in the LS is collected by PMTs. These PMTs are mounted on PMT modules, which are fixed on the stainless steel (SS) structure supporting the CD. Among the LPMTs, 5,000 are dynode PMTs manufactured by Hamamatsu Photonics K.K. (HPK), while the remaining LPMTs are a new type of Micro-Channel Plate Photomultiplier Tubes (MCP-PMTs) manufactured by Northern Night Vision Technology Co. (NNVT). Furthermore, a fraction of the NNVT MCP-PMTs, manufactured with improved technologies during photocathode fabrication, are labeled as NNVT HQE MCP-PMT. The arrangement of these LPMTs has been extensively studied, and their photocathode coverage has been determined to be 75%. Each LPMT is equipped with a protection cover to prevent chain explosions in case of PMT breakage under the high water pressure. The protection consists of a front acrylic cover with a thickness of approximate 10 mm that follows the shape of the photocathode with a minimum 2 mm gap filled with water, working together with a back SS cover to provide full protection [23]. SPMTs are mounted in the gap between LPMTs, providing an additional $\sim$3photocathode coverage. The PMT modules made of SS play a role of optical isolation between the outer and inner WP. In addition, a comprehensive calibration system is also implemented in the CD to control systematic errors in the energy scale and energy response of the detector [24].
The output signals from each LPMT are duplicated and amplified by two Trans-Impedance Amplifiers (TIA) with different amplification factors. This configuration allows for a good signal-to-noise ratio for low-energy events, such as the inverse beta decay process (IBD), with the high gain configuration. At the same time, it maintains a large dynamic range with the low gain configuration to detect high-energy events. The amplified analog signals from each TIA are then digitized using a custom flash analog-to-digital converter (FADC). The FADC provides a sampling rate of 1 GHz and a resolution of 14 bits. The digitized waveform data are processed in the Field Programmable Gate Array (FPGA) to extract the charge and time information. The charge and time values of triggered events, along with the waveform data, are then sent to the Data Acquisition (DAQ) cluster. In the DAQ cluster, complex online event classification algorithms are employed to decide whether to save the full waveform or just the charge and time information onto the disk. This decision is made in order to optimize the data bandwidth and storage capacity. For IBD events, the waveform data is typically recorded to allow for more precise processing and analysis offline. More details of the LPMT electronics and trigger design are available in [25, 26, 27]. Regarding the SPMT electronics, CatiROC chips are used to readout the signals and only charge and time information are saved. More information is outlined in [28].
@@
Figure 1: Schematic drawing of the JUNO detector.
@@ 3 Origin of the energy resolution
Reactor electron antineutrinos with energy exceeding 1.8 MeV can be detected within the LS target from interactions of the products of the IBD process: $\bar{\nu}_{e}+p\to e^{+}+n$. The positron ($e^{+}$), being much lighter than the neutron ($n$), carries away almost all the kinetic energy of the electron antineutrino.It quickly deposits its kinetic energy in the LS and annihilates into two 0.511 MeV gammas, generating a prompt signal within an energy range of 1 to 12 MeV. After scattering in the LS with an average lifetime of approximately $200~{}\mu$s, the neutron is captured by a hydrogen or a carbon nucleus, consequently producing a delayed gamma signal of 2.22 MeV or 4.95 MeV, respectively. The coincidence signature of the prompt and delayed signals can effectively distinguish the IBD signals from the backgrounds. An accurate measurement of the positron energy and understanding its energy resolution in the JUNO detector are crucial for the NMO determination. The impact from the neutron recoils will not be discussed in this article, but is included in the NMO paper [14].
The energy deposition ($E_{\mathrm{dep}}$) of positrons primarily occurs through excitation and ionization processes. Since the concentrations of PPO and bis-MSB in the LS are relatively low, the deposited energy is mainly transferred to LAB molecules. Subsequently, the excited LAB molecules can transfer their energy to fluor molecules through complex molecular-scale processes [30]. During the de-excitation of the fluor molecules, scintillation photons are emitted. However, the number of emitted scintillation photons is not linearly proportional to the deposited energy due to the quenching phenomenon, where some excited molecules release energy without radiation emission. Various models exist to describe this quenching effect, the most common one is the semi-empirical formula proposed by Birks [31], which is used in this study,
@@
where $N_{\textrm{ScintOP}}$ represents the number of optical photons produced by scintillation process, $Y$ corresponds to the scintillation light yield in unit of photons per MeV without the presence of quenching, $dE/dx$ denotes the energy loss per unit path length (also known as the stopping power), and $kB$ is the Birks’ coefficient that depends on the particle type. Given a fixed energy loss, $N_{\textrm{ScintOP}}$ follows a Poisson distribution due to the random nature of the molecular-scale energy transfer process of scintillation, where the variance of the distribution is determined by $Y$. However, the original Poisson distribution is distorted by fluctuations in the energy loss described by a straggling function, the quenching effect, and the following random secondary particle generation processes:
@@
  • Gammas can be generated through Bremsstrahlung radiation and positron annihilation, and interact through photoelectric effect, Compton scattering, and pair production. Then the secondary electrons and positrons deposit their energies and produce scintillation photons through the aforementioned processes.
When the positron and its charged secondaries move through the LS at a velocity higher than the group velocity of light in the LS, Cherenkov photons are emitted. The Frank-Tamm formula [32] is commonly used to calculate the number of Cherenkov photons produced per wavelength per unit path length travelled by particles of charge $ze$,
@@
In the formula, $\alpha$ represents the fine structure constant, $\beta$ is the ratio of the particle’s velocity to the speed of light in vacuum, and $n$ is the wavelength-dependent refractive index (RI) of the LS. According to Eq. (2), the number of Cherenkov photons is inversely proportional to the square of the wavelength ($\lambda$), resulting in a higher Cherenkov light yield at shorter wavelengths. However, there is very limited information on the RI at short wavelengths, which introduces significant uncertainties in estimating $N_{\textrm{CherenOP}}$. Although the Cherenkov process contributes additional light, it has a detrimental effect on the energy resolution. This is because the Cherenkov light yield depends on the path length of the charged particle in the LS, which can vary significantly due to the generation of secondary particles. This variation in path length can cause the Cherenkov photon distribution to deviate from Poisson statistics, leading to a degradation of the energy resolution.
Scintillation and Cherenkov photons produced in the LS may undergo various processes as they propagate through the detector, including absorption, scattering, reflection, and transmission at material boundaries. Absorption within the LS can be modeled as a competition among the LAB, PPO, and bis-MSB molecules. The absorption length of each component determines the probability of photon absorption. When photons are absorbed, they may be re-emitted again, and this re-emission probability is determined by the quantum yield (QY) of the fluors in the LS. Scattering within the LS is dominated by Rayleigh scattering rather than Mie scattering, because LS is expected to have an extremely low concentration of large-sized impurities. The probability of Rayleigh scattering depends on the scattering length of the photons. The propagation of optical photons inside the LS can be well modeled with the above optical parameters provided. Only a fraction of these photons will be detected by the PMTs and converted into PE. The total number of PE ($N_{\textrm{PE}}$) is influenced by factors such as the scintillation light yield, the quenching effect, the LS refractive index, the light collection efficiency during transportation, and the photon detection efficiency (PDE) of the PMTs. A simple energy reconstruction method is to divide $N_{\textrm{PE}}$, including contributions from both scintillation PE ($N_{\textrm{ScintPE}}$) and Cherenkov PE ($N_{\textrm{CherenPE}}$), by the energy scale which is defined as the average PE number per MeV calibrated by 2.2 MeV gammas from neutron capture on hydrogen at the center of CD. The number of PE produced in both scintillation and Cherenkov processes are not proportional to $E_{\textrm{dep}}$, leading to LS non-linearity (LSNL) denoted as:
@@
where $E_{\textrm{vis}}$ is the visible energy defined as the expected reconstructed energy assuming a perfect energy resolution. Due to the aforementioned fluctuations in the energy loss and Cherenkov process, $N_{\textrm{PE}}$ does not follow a simple Poisson distribution with the variance of $N_{\textrm{PE}}$ and exhibits a larger standard deviation. To account for these fluctuations, a more general formula can be introduced as follows:
@@
$\sigma_{\textrm{PE}}=\sqrt{\sigma_{\textrm{ScintPE}}^{2}+\sigma_{\textrm{CherenPE}}^{2}+2\cdot cov[N_{\textrm{ScintPE}},N_{\textrm{CherenPE}}]}>\sqrt{N_{\textrm{PE}}},$ (4)
@@
Equation (4) includes the correlation between the fluctuations in the scintillation and Cherenkov radiation which has been observed in the JUNO simulation and will be further discussed in the following section. The value of $N_{\textrm{PE}}$ also depends on the position of the IBD event. This position-dependent effect is referred to as the detector non-uniformity with respect to the amount of the collected photons for the same energy release, which arises from geometrical factors. Consequently, the energy resolution is also position-dependent in the JUNO detector.
@@
The PMT and electronics responses can introduce additional fluctuations to the detected $N_{\textrm{PE}}$. The PMT response includes various factors such as the single photoelectron (SPE) resolution, which arises from the inherent fluctuations in the dynode or MCP charge amplification process. Additionally, effects such as dark count rate (DCR) and afterpulses can contribute to the fluctuations if they occur within the signal readout window and mix with the PE signal. The transit time and transit time spread (TTS) of the PMTs can impact the accuracy of vertex reconstruction and further influence the energy resolution, due to the detector non-uniformity. On the electronics side, the presence of electronics noise can further smear the charge measurements. The digitization process can also contribute to the overall fluctuations. All of these factors and parameters have been well characterized during the PMT mass testing [17] and electronics development phases [33]. Their impacts on the resolution depend also on the specific vertex and energy reconstruction algorithms employed.
@@
Based on the waveform recorded in each readout channel, the charge and time information can be extracted by the waveform reconstruction algorithm. Then, the charge can be calibrated and converted to the number of PE using PMT calibration algorithms. Eventually, the event position and energy can be obtained from the vertex and energy reconstruction algorithms using charge and time information of each readout channel as inputs. During reconstruction, additional fluctuations, from PMT calibration, residual energy non-uniformity, among others, will be added to the reconstructed energy, where their amplitudes depend on the performance of the algorithms.
@@
The main factors that may impact the energy resolution in the JUNO detector are summarized below, and also illustrated in Fig. 2 with key factors indicated:
@@
    Scintillation light yield and ionization quenching effect during the scintillation photons production.
@@
All of the mentioned items are carefully considered in the full JUNO detector simulation chain and algorithms of event reconstruction. Items 1-4, which include the energy deposition, scintillation (Sec. 4.2) and Cherenkov light production (Sec. 4.3), photon propagation (Sec. 4.4) and detection (Sec. 4.6), are taken into account in the simulation. Item 5, which involves the PMT and electronics responses (Sec. 5.1 and Sec. 5.2), is modeled in a dedicated electronics simulation. Item 6 is included in the data processing of the calibration (Sec. 6.1) and reconstruction (Sec. 6.2).
Figure 2: Summary of the key factors impacting the energy resolution throughout the processes of light production, propagation, detection, calibration and reconstruction.

### v2

@@ Abstract
This paper presents an energy resolution study of the JUNO experiment, incorporating the latest knowledge acquired during the detector construction phase. The determination of neutrino mass ordering in JUNO requires an exceptional energy resolution better than 3% at 1 MeV. To achieve this ambitious goal, significant efforts have been undertaken in the design and production of the key components of the JUNO detector. Various factors affecting the detection of inverse beta decay signals have an impact on the energy resolution, extending beyond the statistical fluctuations of the detected number of photons, such as the properties of the liquid scintillator, performance of photomultiplier tubes, and the energy reconstruction algorithm. To account for these effects, a full JUNO simulation and reconstruction approach is employed. This enables the modeling of all relevant effects and the evaluation of associated inputs to accurately estimate the energy resolution. The results of study reveal an energy resolution of 2.95% at 1 MeV. Furthermore, this study assesses the contribution of major effects to the overall energy resolution budget. This analysis serves as a reference for interpreting future measurements of energy resolution during JUNO data collection. Moreover, it provides a guideline for comprehending the energy resolution characteristics of liquid scintillator-based detectors.
@@ 1 Introduction
The discovery of neutrino oscillations has established that at least two neutrino mass eigenstates are massive. This unambiguous conclusion has provided the first evidence of new physics beyond the Standard Model, which assumes neutrinos are massless. In the framework of three neutrino oscillations, six fundamental parameters are involved to describe the oscillation phenomena. Solar and reactor neutrino experiments have determined the mass-squared difference $\Delta m^{2}_{21}$ and mixing angle $\sin^{2}\theta_{12}$. Atmospheric and accelerator experiments have measured $|\Delta m^{2}_{32}|$ (or $|\Delta m^{2}_{31}|$) and $\sin^{2}\theta_{23}$. Reactor antineutrino experiments have provided the most precise determination of $\sin^{2}\theta_{13}$ and have also achieved a comparable precision of $|\Delta m^{2}_{32}|$ with the measurements from atmospheric and accelerator experiments. The two independent mass-squared differences and the three mixing angles have been measured with a precision of a few percent [1]. However, the sign of $\Delta m^{2}_{32}$ (referred to as neutrino mass ordering; NMO) and the value of the CP violating phase $\delta_{\textrm{CP}}$ remain partially unknown [2, 3, 4]; thus, several experiments [5, 6, 7, 8] have been proposed to determine these two crucial parameters.
@@
Over the past few years several factors related to the NMO sensitivity in JUNO have undergone changes. As a result, the JUNO collaboration recently updated the NMO sensitivity using more realistic inputs [18]. These updates include the incorporation of updated reactor thermal powers and new constraints on the reactor neutrino spectrum that we expect to obtain from the TAO [19] satellite detector. Additionally, the accidental backgrounds have been revised based on radioassay results of the detector’s raw materials and the estimation of radioactive impurities during the detector installation [20]. The cosmogenic backgrounds have been re-evaluated due to changes in the overburden, and new backgrounds from global reactors and atmospheric neutrinos have been included. Furthermore, the detector energy resolution has been updated through full detector simulation and event reconstruction, which take into account a better understanding of the detector structure, as well as the performance of the LS and photomultiplier tubes (PMTs) [21, 22]. This paper serves as a complementary report to the updated NMO sensitivity paper of JUNO and provides a detailed description of the latest understanding regarding the energy resolution prediction in the JUNO experiment, which is essential for NMO analysis.
@@
Previous LS-based neutrino detectors, such as the Borexino [23] and KamLAND [24] experiments, have achieved energy resolution levels of 5% and 6.5% at 1 MeV, respectively. In these experiments, statistical fluctuations dominate the energy resolution, with reported photoelectron (PE) yields of 511 PE/MeV and 250 PE/MeV, respectively. The JUNO detector was designed with a focus on achieving high LS light yield, high transparency, and high photon collection efficiency to minimize the statistical fluctuation of the detected PE. The LS recipe has been optimized to maximize the light yield using one of the antineutrino detectors in the Daya Bay experiment [25]. Additionally, 17,612 high-detection-efficiency 20-inch PMTs (LPMTs) [21] and 25,600 3-inch PMTs (SPMTs) [26] have been deployed, resulting in a total photocathode coverage of 78%. However, due to the unprecedented energy resolution requirement of JUNO, several other factors, which were usually negligible in previous experiments, significantly contribute to the energy resolution and must be carefully evaluated. These factors include the quenching effect in LS, Cherenkov radiation, electronics, and PMT responses. All of these effects have been modeled in the full detector simulation and will be thoroughly discussed in this paper, along with the impacts from the vertex and energy reconstruction algorithms.
@@
The remainder of this article is organized as follows. We first provide an introduction to the JUNO detector in Sec. 2, followed by a summary of the main factors that impact the energy resolution throughout the processes of light production, propagation, and detection in Sec. 3. The full detector simulation is then presented in Sec. 4, focusing on the optical models of the LS and PMTs. Updates and improvements made to these models based on the latest knowledge and measurements are highlighted. The approaches used to determine key parameters in these models are also discussed. Additionally, the electronics simulation is introduced in Sec. 5, which provides a comprehensive model of the detector response. Section 6 describes the calibration and event reconstruction procedures and reports the newly predicted energy resolution of JUNO, leading to a decomposition of the energy resolution to identify the major contributing factors in Sec. 7 . Finally, a summary is given in Sec. 8.
@@ 2 The JUNO detector
A schematic of the JUNO detector is shown in Fig. 1. The detector consists of three main components: the central detector (CD), water pool (WP), and top tracker (TT).
@@
Scintillation and Cherenkov light produced in the LS are collected by PMTs. PMT modules are the basic units used to assemble PMTs on the stainless steel (SS) supporting structure of the CD, accommodating multiple PMTs in each module. Among the LPMTs, 5,000 are dynode PMTs manufactured by Hamamatsu Photonics K.K. (HPK), while the remaining LPMTs are a new type of Micro-Channel Plate Photomultiplier Tubes (MCP-PMTs) [28] manufactured by Northern Night Vision Technology Co. (NNVT). Furthermore, a fraction of the NNVT MCP-PMTs, manufactured with improved technologies during photocathode fabrication, are labeled as NNVT HQE MCP-PMT. The arrangement of these LPMTs has been extensively studied, and their photocathode coverage has been determined to be 75%. Each LPMT is equipped with a protection cover to prevent chain explosions in case of PMT breakage under the high water pressure. The protection consists of a front acrylic cover with a thickness of approximately 10 mm that follows the shape of the photocathode with a minimum 2 mm gap filled with water, working together with a back SS cover to provide full protection [29]. SPMTs are mounted in the gap between LPMTs, providing an additional $\sim$3photocathode coverage. The PMT modules made of SS provide optical isolation between the outer and inner WP. In addition, a comprehensive calibration system is implemented in the CD to control systematic errors in the energy scale and energy response of the detector [30].
The output signals from each LPMT are duplicated and amplified by two trans-impedance amplifiers (TIA) with different amplification factors. This configuration allows for a good signal-to-noise ratio for low-energy events, such as the inverse beta decay (IBD) process, with the high gain configuration. At the same time, it maintains a large dynamic range with the low gain configuration to detect high-energy events. The amplified analog signals from each TIA are then digitized using a custom flash analog-to-digital converter (FADC). The FADC provides a sampling rate of 1 GHz and resolution of 14 bits. The digitized waveform data are processed in the field programmable gate array (FPGA) to extract the charge and time information. The charge and time values of triggered events, along with the waveform data, are then sent to the data acquisition (DAQ) cluster. In the DAQ cluster, complex online event classification algorithms are employed to decide whether to save the full waveform or just the charge and time information onto the disk. This decision is made to optimize the data bandwidth and storage capacity. For IBD events, the waveform data are typically recorded to allow for more precise processing and analysis offline. More details of the LPMT electronics and trigger design are available in [31, 32, 33]. Regarding the SPMT electronics, CatiROC chips are used to readout the signals, and only charge and time information are saved. More information is outlined in [34].
@@
Figure 1: Schematic of the JUNO detector. TT: top tracker, LS: liquid scintillator, CD: central detector, SS: stainless steel.
@@ 3 Origin of the energy resolution
Reactor electron antineutrinos with energy exceeding 1.8 MeV can be detected within the LS target from interactions of the products of the IBD process: $\bar{\nu}_{e}+p\to e^{+}+n$. The positron ($e^{+}$), being much lighter than the neutron ($n$), carries away almost all the kinetic energy of the electron antineutrino. It quickly deposits its kinetic energy in the LS and annihilates into two 0.511 MeV gammas, generating a prompt signal within an energy range of 1 to 12 MeV. After scattering in the LS with an average lifetime of approximately $200~{}\mu$s, the neutron is captured by a hydrogen or carbon nucleus, consequently producing a delayed gamma signal of 2.22 or 4.95 MeV, respectively. The coincidence signature of the prompt and delayed signals can effectively distinguish the IBD signals from the backgrounds. An accurate measurement of the positron energy and understanding of its energy resolution in the JUNO detector are crucial for NMO determination. The impact from the neutron recoils will not be discussed in this article, but it is considered in the NMO paper [18].
The energy deposition ($E_{\mathrm{dep}}$) of positrons primarily occurs through excitation and ionization processes. Because the concentrations of PPO and bis-MSB in the LS are relatively low, the deposited energy is mainly transferred to LAB molecules. Subsequently, the excited LAB molecules can transfer their energy to fluor molecules through complex molecular-scale processes [36]. During the de-excitation of the fluor molecules, scintillation photons are emitted. However, the number of emitted scintillation photons is not proportional to the deposited energy due to the quenching phenomenon [37], where some excited molecules release energy without radiation emission. Various models exist to describe this quenching effect; the most common one is the semi-empirical formula proposed by Birks [38], which is used in this study:
@@
where $N_{\textrm{ScintOP}}$ represents the number of optical photons produced by scintillation process, $Y$ corresponds to the scintillation light yield in units of photons per MeV without the presence of quenching, $dE/dx$ denotes the energy loss per unit path length (also known as the stopping power), and $kB$ is the Birks coefficient, which depends on the particle type. Given a fixed energy loss ($dE$) over a short path length of $dx$, $dN_{\textrm{ScintOP}}$ follows a Poisson distribution due to the random nature of the molecular-scale energy transfer process of scintillation, where the variance of the distribution is determined by $Y$. However, the total number of scintillation photons ($N_{\textrm{ScintOP}}$) produced by a particle no longer follows the Poisson distribution due to fluctuations in the energy loss described by a Landau distribution, the quenching effect modeled by Birks’ Law, and the following random secondary particle generation processes:
@@
  • Gammas can be generated through Bremsstrahlung radiation and positron annihilation, and they interact through photoelectric effect, Compton scattering, and pair production. Then, the secondary electrons and positrons deposit their energies and produce scintillation photons through the aforementioned processes.
When the positron and its charged secondaries move through the LS at a velocity higher than the group velocity of light in the LS, Cherenkov photons are emitted. The Frank-Tamm formula [39] is commonly used to calculate the number of Cherenkov photons produced per wavelength per unit path length travelled by particles of charge $ze$:
@@
In this formula, $\alpha$ represents the fine structure constant, $\beta$ is the ratio of the particle’s velocity to the speed of light in vacuum, and $n$ is the wavelength-dependent refractive index of the LS. According to Eq. (2), the number of Cherenkov photons is inversely proportional to the square of the wavelength ($\lambda$), resulting in a higher Cherenkov light yield at shorter wavelengths. However, there is very limited information on the refractive index at short wavelengths, which introduces significant uncertainties in estimating $N_{\textrm{CherenOP}}$. Although the Cherenkov process contributes additional light, it has a detrimental effect on the energy resolution. This is because the Cherenkov light yield depends on the path length of the charged particle in the LS, which can vary significantly due to the generation of secondary particles. This variation in path length can cause the Cherenkov photon distribution to deviate from Poisson statistics, leading to a degradation of the energy resolution.
Scintillation and Cherenkov photons produced in the LS may undergo various processes as they propagate through the detector, including absorption, scattering, reflection, and transmission at material boundaries. Absorption within the LS can be modeled as a competition among the LAB, PPO, and bis-MSB molecules. The absorption length of each component determines the probability of photon absorption. When photons are absorbed, they may be re-emitted again, and this re-emission probability is determined by the quantum yield (QY) of the fluors in the LS. Scattering within the LS is dominated by Rayleigh scattering rather than Mie scattering because LS is expected to have an extremely low concentration of large-sized impurities. The probability of Rayleigh scattering depends on the scattering length of the photons. The propagation of optical photons inside the LS can be well modeled with optical parameters such as absorption length, re-emission probability, Rayleigh scattering length, and refractive indices of the detector components. Only a fraction of these photons will be detected by the PMTs and converted into PE. The total PE number ($N_{\textrm{PE}}$) is influenced by factors such as the scintillation light yield, quenching effect, LS refractive index, light collection efficiency during transportation, and photon detection efficiency (PDE) of the PMTs. A simple energy reconstruction method is to divide $N_{\textrm{PE}}$, including contributions from both scintillation PE ($N_{\textrm{ScintPE}}$) and Cherenkov PE ($N_{\textrm{CherenPE}}$), by the energy scale, which is defined as the average PE number per MeV calibrated by 2.2 MeV gammas from neutron capture on hydrogen at the center of the CD. The average PE number produced in both scintillation and Cherenkov processes is not proportional to $E_{\textrm{dep}}$, leading to LS non-linearity (LSNL), denoted as
@@
where $E_{\textrm{vis}}$ is the visible energy defined as the expected reconstructed energy assuming a perfect energy resolution. Due to the aforementioned fluctuations in the energy loss and Cherenkov process, $N_{\textrm{PE}}$ does not follow a simple Poisson distribution but exhibits a larger standard deviation. To account for these fluctuations, a more general formula can be introduced as follows:
@@
$\sigma_{\textrm{PE}}=\sqrt{\sigma_{\textrm{ScintPE}}^{2}+\sigma_{\textrm{CherenPE}}^{2}+2\cdot cov[N_{\textrm{ScintPE}},N_{\textrm{CherenPE}}]}>\sqrt{\langle N_{\textrm{PE}}\rangle},$ (4)
@@
where $\sigma_{\textrm{PE}}$ is the standard deviation of the $N_{\textrm{PE}}$ distribution. Equation (4) includes the correlation between the fluctuations in the scintillation and Cherenkov radiation which has been observed in the JUNO simulation and will be further discussed in the following section. The value of $N_{\textrm{PE}}$ also depends on the position of the IBD event. This position-dependent effect is referred to as the detector non-uniformity with respect to the amount of collected photons for the same energy release, which arises from geometrical factors. Consequently, the energy resolution in the JUNO detector is also position-dependent.
@@
The PMT and electronics responses can introduce additional fluctuations to the detected $\langle N_{\textrm{PE}}\rangle$. The PMT response includes various factors, such as the single photoelectron (SPE) resolution, which arises from the inherent fluctuations in the dynode or MCP charge amplification process. Additionally, effects such as dark count rate (DCR) and afterpulses can contribute to the fluctuations if they occur within the signal readout window and mix with the PE signal. The transit time and transit time spread (TTS) of the PMTs can impact the accuracy of vertex reconstruction and further influence the energy resolution due to the detector non-uniformity. On the electronics side, the presence of electronics noise can further smear the charge measurements. The digitization process can also contribute to the overall fluctuations. All of these factors and parameters have been well characterized during the PMT mass testing [21] and electronics development phases [40]. Their impacts on the resolution depend also on the specific vertex and energy reconstruction algorithms employed.
@@
Based on the waveform recorded in each readout channel, the charge and time information can be extracted by the waveform reconstruction algorithm. Then, the charge can be calibrated and converted to the PE number using PMT calibration algorithms. Eventually, the event position and energy can be obtained from the vertex and energy reconstruction algorithms using charge and time information of each readout channel as inputs. During reconstruction, additional fluctuations, PMT calibration, residual energy non-uniformity, and other factors are added to the reconstructed energy, where their amplitudes depend on the performance of the algorithms.
@@
The main factors that may impact the energy resolution in the JUNO detector are summarized as follows (also illustrated in Fig. 2):
@@
    Scintillation light yield and ionization quenching effect during scintillation photon production.
@@
All of the above items are carefully considered in the full JUNO detector simulation chain and algorithms of event reconstruction. Items 1-4, which include the energy deposition, scintillation (Sec. 4.2) and Cherenkov light production (Sec. 4.3), photon propagation (Sec. 4.4), and detection (Sec. 4.6), are taken into account in the simulation. Item 5, which involves the PMT and electronics responses (Sec. 5.1 and Sec. 5.2), is modeled in a dedicated electronics simulation. Item 6 is included in the data processing of the calibration (Sec. 6.1) and reconstruction (Sec. 6.2).
Figure 2: Summary of the key factors impacting the energy resolution throughout the processes of light production, propagation, detection, calibration, and reconstruction.